A total of $7000 is deposited into two simple interest accounts. In one account the annual simple interest rate is 3%, and in the second account the annual simple interest rate is 7%. The amount of interest earned for 1 year was $250. How much was invested in each account?
Answer by math_tutor2020(3817)


The simple interest formula is
i = P*r*t
where,
i = interest
P = deposit
r = interest rate in decimal form
t = number of years

Let
x = amount deposited at 3% rate
7000-x = amount deposited at 7% rate
The two items add to $7000

For the 3% account
i = P*r*t
i = x*0.03*1
i = 0.03x

For the 7% account
i = P*r*t
i = (7000-x)*0.07*1
i = 490-0.07x

Total interest = 0.03x+(490-0.07x)
Total interest = -0.04x+490

Set that equal to $250 and solve for x.
Total interest = 250
-0.04x+490 = 250
-0.04x = 250-490
-0.04x = -240
x = -240/(-0.04)
x = 6000
$6000 was invested at 3% interest rate.

Then use mental math to determine the remaining amount invested must be $1000

Or you can say: 7000-x = 7000-6000 = 1000

Check:
6000+1000 = 7000
6000*0.03+1000*0.07 = 250
Both conditions are verified.


Answer:
$6000 invested at 3%
$1000 invested at 7%